PLEASE SHOW ALL TABLES USED!!!!! AND HAVE NEAT HANDWRITING!!

A steam generator discharges water vapor at 10 MPa and 640 degrees Celsius. The first turbine stage expands the vapor to 2 MPa, which is extracted to a feedwater heater. The remaining portion expands through the second turbine stage to a condenser pressure of 10 kPa. The heat transfer into the steam generator is 5 MW, and the first pump operates at 60% isentropic efficiency. The enthalpies exiting the feedwater heater and the second pump are given as 909 kJ/kg and 918 kJ/kg, respectively. Determine

A. Mass flow rate exiting the first pump (kg/s)

B. Total pump work (kW)

C. Thermal efficiency
Answer #1

Iven 0-Table A-4, Ar5, A- G P10M Pa 2 h,-372364 KIJK. O1z 701354 KT|K9- 3173.5 Klkg SI=S2 OMP るC10KPへ2。=21-8 KI|kI 2MP S = SThe flou rate eriting The first is, m - mmass PLLP Gen =5000K W = mi(hi-ha) S000 (3723.64-918) kgls mm1-0.239) I-18) m135kgls